It takes you all of 0.018s to initially touch and then catch a 0.600kg football travelling at 16.0m/s

a. what is the change in momentum for the football?
b. what is the impulse?
c. what is the force that must be exerted to stop the ball?
1
Expert's answer
2010-11-10T02:25:23-0500
Change in the momentum: Δp = mΔv = 0.6*16 = 9.6 [kg* m/s]

The impulse equals to momentum change: I = Δp = 9.6 [kg* m/s]

The force:
F = mΔv/t = 0.600*16/0.018 = 533.33 [N]